HighGrowth Company has a stock price of

$ 20$20.

The firm will pay a dividend next year of $1.00​, and its dividend is expected to grow at a rate of 4.0% per year thereafter. What is your estimate of​ HighGrowth's cost of equity​ capital?

The required return​ (cost of​ capital) of levered equity is _______%.​ (Round to one decimal​ place.)
Answer #1

Ans 9.0%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
20 = 1 / (Ke - 4%)
Ke - 4% = 1 / 20
Ke - 4% = 5.00%
Ke = 9.0%